Use Z substitution
I= ∫ dx /(1+sinx)
1
Expert's answer
2020-09-28T16:43:49-0400

"\displaystyle\textsf{Let}\hspace{0.1cm}I = \int \frac{1}{1 + \sin{x}}\hspace{0.1cm} \mathrm{d}x \\

\textsf{Substitute}\hspace{0.1cm} Z = \tan\left(\frac{x}{2}\right) \\

\begin{aligned}
I &= \int \frac{1}{1 + \frac{2Z}{1 + Z^2}}\hspace{0.1cm} \frac{2\mathrm{d}Z}{1 + Z^2}\\
&= \int \frac{1}{1 + Z^2 + 2Z}\hspace{0.1cm}2\mathrm{d}Z = \int \frac{2}{(Z + 1)^2}\hspace{0.1cm}\mathrm{d}Z \\&= \frac{-2}{1 + Z} + C \\&= \frac{-2}{1 + \tan\left(\frac{x}{2}\right)} + C\\
\end{aligned}\\

\textsf{Multiplying the numerator and}\\ \textsf{denominator} \hspace{0.1cm}\textsf{of}\hspace{0.1cm} I\hspace{0.1cm}\textsf{by} \hspace{0.1cm}\cos\left(\frac{x}{2}\right)\sin\left(\frac{x}{2}\right), \hspace{0.1cm} \textsf{we have;}\\
I =\frac{-2\cos\left(\frac{x}{2}\right)\sin\left(\frac{x}{2}\right)}{\cos\left(\frac{x}{2}\right)\sin\left(\frac{x}{2}\right) + \sin^2\left(\frac{x}{2}\right)} + C\\

\textsf{It is also known that}\hspace{0.1cm} \\\sin{(x)} = 2\cos\left(\frac{x}{2}\right)\sin\left(\frac{x}{2}\right) \hspace{0.1cm}\&\hspace{0.1cm}\\\cos{(x)} = 1 - 2\sin^2\left(\frac{x}{2}\right)\\

\begin{aligned}
\therefore I &=\frac{-2\cos\left(\frac{x}{2}\right)\sin\left(\frac{x}{2}\right)}{\cos\left(\frac{x}{2}\right)\sin\left(\frac{x}{2}\right) + \sin^2\left(\frac{x}{2}\right)} + C\\
&=\frac{-\sin{x}}{\frac{\sin{x}}{2} + \frac{1 - \cos{x}}{2}} + C= \frac{-2\sin{x}}{1 - \cos{x} + \sin{x}} + C\\&=\frac{2}{\cot{x} -\cosec{x} - 1} + C
\end{aligned}\\


\therefore I = \int \frac{1}{1 + \sin{x}}\hspace{0.1cm} \mathrm{d}x = \frac{2}{\cot{x} -\cosec{x} - 1} + C"